I want to get objective short tricks in aieee and iit which can save my time mainly in maths.you also can tell me the site.i want to negative marking in aieee and iit. ialso have some questions pending.please check and send.

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear vipul

There are many tricks to solve tough problem

but a single trick u can't use for every problem

tricks changes accroding to question that u can learn by solving question

more question u will solve u will learn more tricks u will develope ur own..

for example

              if (a+x+x2)6n= a0 +a1x+a2x2+........

                  then find a0+a3+.......  ?

Sol:

 here if u will solve it will take lot of time

but u can take n=1/3 then u can find a0+a3+.......

when you put n=1/3  

 (a+x+x2)6n =(a+x+x2)

expand this  and compair with R H S term

 


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

 

OCEAN GODRE
13 years ago

short tricks can only and only be developed through practice of a large no of questions..thr is no particular books giving short tricks..